What would be the effect on the volume of an ideal gas if the pressure is doubled and the temperature in kelvins is halved? Defend this response.
Answer #1

use combined gas law:

Pi*Vi/Ti = Pf*Vf/Tf

Given:

Pf = 2*Pi

Tf = Ti/2

use:

Pi*Vi/Ti = Pf*Vf/Tf

Pi*Vi/Ti = 2*Pi*Vf/(Ti/2)

Pi*Vi/Ti = 4*Pi*Vf/Ti

Vi = 4Vf

Vf=Vi/4

Answer:

Volume becomes 1/4 times the initial value